How much would the equivalent

resistance in a parallel circuit be if it had a 3Ω, a

6 Ω, a 5 Ω, and a 2 Ω resistor?


1
Expert's answer
2021-04-13T06:39:58-0400

"R_1=3\Omega\\R_2=6\Omega\\R_3=5\Omega\\R_4=2\Omega"

"\\\Rightarrow\dfrac{1}{R_{eq}}=\dfrac{1}{R_1}+\dfrac{1}{R_2}+\dfrac{1}{R_3}+\dfrac{1}{R_4}"

"\Rightarrow\dfrac{1}{R_{eq}}=\dfrac{1}{3}+\dfrac{1}{6}+\dfrac{1}{5}+\dfrac{1}{2}"

"\Rightarrow\dfrac{1}{R_{eq}}=\dfrac{10+5+6+15}{30}"

"\Rightarrow R_{eq}=\dfrac{30}{36}=0.833\space\Omega"

Req = 0.833 "\Omega"